OpenStudy (anonymous):

Solve by factoring: 8x^2=24x

OpenStudy (turingtest):

8x^2-24x=0 8x(x-3)=0 x={0,3} ask if you need explanation

OpenStudy (anonymous):

please do

OpenStudy (anonymous):

???

OpenStudy (anonymous):

like how did you get 8x(x-3)=0

OpenStudy (turingtest):

ok, first we get everything on one side with the other side equal to zero: 8x^2=24x 8x^2-24x=0 Now factor out the 8x. Notice that everything here is divisible by 8x (that's what we mean by "factor out the 8x) 8x(x-3) if this step is giving you trouble, think about at least factoring out the x: 8x^2-24x=0 x(8x-24)=0 since either x or what is within the parentheses MUST be zero, we can solve these two expressions seperately: x=0 8x-24=0 x=0 8x=24 x=0 x=3 so whether you factor out the 8 along with the x or not, you see you get the same answer x={0,3}

OpenStudy (anonymous):

\[8x^2=24x\] Subtracting 24x from both sides we will get : \[8x^2-24x=0\] Do you noticed that this is a quadratic equation? However, now you could factor out the 8x: \[8x(x-3)=0 \Rightarrow x =0 \text{ or } x =3\]